OpenStudy (cassieforlife5):

I'm totally confused on how to determine if a series is convergent or divergent. So for this series: -20+100-500 I believe that it is convergent because r=-5. However, I'm supposed to use the comparison test on this question: (1/49)+(1/64)+(1/81)+... And I have no idea how to go about it. Could you also help me understand how to find the general term of the series? Any help would be greatly appreciated!

OpenStudy (anonymous):

The first series (assuming infinite series), \[-20+100-500+\cdots=4\sum_{k=0}^\infty(-5)^{k+1}\] diverges precisely because the ratio is \(r=-5\). A geometric series can only converge if \(|r|<1\).

OpenStudy (anonymous):

For the second series, what you have is the sum of reciprocal squares, \[\frac{1}{49}+\frac{1}{64}+\frac{1}{81}+\cdots=\frac{1}{7^2}+\frac{1}{8^2}+\frac{1}{9^2}+\cdots=\sum_{k=7}^\infty\frac{1}{k^2}\] This series converges because it's smaller than another almost identical series, \[\sum_{k=1}^\infty\frac{1}{k^2}=1+\frac{1}{4}+\frac{1}{9}+\cdots\] (the index is different). The "comparison" series can be shown to converge using the integral test. The reason series number 2 converges is because it has fewer terms, so it adds up to a number less than the comparison series' sum.